# Enhancing reliability in prediction intervals using point forecasters: Heteroscedastic Quantile Regression and Width-Adaptive Conformal Inference

### v1

@@ 3.2 Width-Adaptive Conformal Inference
The other property we are looking for is to have the same level of confidence regardless of the difficulty of the prediction. That is, the coverage should not depend on the complexity of the situation. It has been established that an indicator of the difficulty is the length of the interval, which in our case is constructed by a quantile regression process. We denote this interval at time $T+1$ by $C_{\alpha,T+1}$. To guarantee the desired property, this interval will be modified by constructing a second one, denoted by $C^{c}_{\alpha,T+1}$, such that it is satisfied that
$\mathbb{P}\left(y_{T+1}\in C^{c}_{\alpha,T+1}|\left|C_{\alpha,T+1}\right|\right)=\mathbb{P}\left(y_{T+1}\in C_{\alpha,T+1}\right)\geq 1-\alpha.$
Since we are trying to achieve a conditional probability, we will draw on ideas from works on conditional Conformal Prediction. In particular, one path to achieve conditional coverage is based on using weighted empirical distributions. That is, not all observations contribute the same. In Tibshirani et al. (2019) it is shown that this weighted based Conformal Prediction also maintains the marginal coverage property. In addition, in works such as Guan (2023) or Han et al. (2022) it is shown that giving weights to the observations as a function of the distance between $X_{i}$ and $X_{T+1}\,\forall i\in\text{Cal}$ results in asymptotic conditional coverage, i.e, if $T\longrightarrow\infty$, then
$\mathbb{P}\left(y_{T+1}\in C_{\alpha,T+1}|X_{T+1}\right)\geq 1-\alpha.$
Intuitively, if one focuses on certain features and considers previous samples with similar values of those features, (asymptotic) conditional coverage with respect to those features is obtained. We can do the same with the length of the interval. As we are working with time series, we will modify the ACI method to apply a different $\alpha$ as a function of time, like the original method, and also as a function of the length of the interval.
Notation 1.
Let $\bm{v}\in\mathbb{R}^{n}$ be a vector of length $n$. The element in position $i$ of $\bm{v},\,i=1,\dots,n$ is denoted by $\bm{v}\left[i\right]$. The $p$ power of $\bm{v}$ is defined as the $p$ power of each one of the elements of $\bm{v}$.
$\bm{v}^{p}=\left(\bm{v}\left[1\right]^{p},\bm{v}\left[2\right]^{p},\dots,\bm{v}\left[n\right]^{p}\right).$
The absolute value of $\bm{v}$ is defined as the absolute values of each one of the elements of $\bm{v}$.
$|\bm{v}|=\left(|\bm{v}\left[1\right]|,|\bm{v}\left[2\right]|,\dots,|\bm{v}\left[n\right]|\right)$
@@
Let $\mathcal{S}$ be the conformity scores (Section 2.2.1). Given a step $h\in\mathbb{R}^{+}$, the 1-d grid long is defined as long = $\left(\text{long}_{\text{min}},\text{long}_{\text{min}}+h,\text{long}_{\text{min}}+2h,\dots,\text{long}_{\text{max}}\right)\in\mathbb{R}^{n}$. Let’s denote by $|C_{\alpha,t}|$ the lenght of the interval of the quantile regession algorithm: $|C_{\alpha,t}|=\hat{u}_{\alpha^{*},t}-\hat{l}_{\alpha^{*},t}\quad\forall\;t\in\text{Cal}\cup\{T+1\}$. Let $\alpha^{*}$ be the objective miscoverage rate. The application of WACI (Width-Adaptive Conformal Inference) over the CQR procedure looks as follows. Let $\bm{\alpha}_{1}=\left(\alpha^{*},\alpha^{*},\dots,\alpha^{*}\right)\in\mathbb{R}^{n},\text{err}_{1}=0,\gamma,\sigma>0.$
@@
$\begin{cases}\bm{\alpha}_{t+1}&=\bm{\alpha}_{t}+\gamma\bm{w}_{t}(\alpha^{*}-\text{err}_{t})\\ \text{err}_{t}&=\begin{cases}1&\text{if }y_{t}\not\in\hat{C}_{\alpha^{*},t}^{c}\\ 0&\text{otherwise}\end{cases}\\ \text{{dist}}_{t}&=|\text{{long}}-|C_{\alpha,t}||\\ \bm{w}_{t}&=\dfrac{\text{exp}\left(\frac{-\text{{dist}}_{t}^{2}}{2\sigma^{2}}\right)}{\max\left\{\text{exp}\left(\frac{-\text{{dist}}_{t}^{2}}{2\sigma^{2}}\right)\right\}}\\ \text{idx}_{t+1}&=\operatorname*{\arg\!\min}\left\{\text{{long}}\left[i\right]-|C_{\alpha,t+1}|\,:\,i=1,\dots,n\right\}\\ \tilde{\alpha}_{t+1}&=\bm{\alpha}_{t+1}\left[\text{idx}_{t+1}\right]\\ C_{\alpha^{*},t+1}^{c}&=\left[\hat{l}_{\alpha^{*}_{t}}^{c},\hat{u}_{\alpha^{*}_{t}}^{c}\right]=\left[\hat{l}_{\alpha^{*},t+1}-Q_{1-\tilde{\alpha}_{t+1}}(\mathcal{S}),\hat{u}_{\alpha^{*},t+1}+Q_{1-\tilde{\alpha}_{t+1}}(\mathcal{S})\right]\\ \end{cases}$ (4)
@@
The first difference that can be seen with the ACI method is that in this case there is not a single $\alpha_{t}$ in each iteration, but a vector $\bm{\alpha}_{t}$. This is done in order to be able to differentiate the real scalar $\tilde{\alpha}_{t}$ that will actually be used in that iteration, which will depend on the length of the interval. That is, each element of the vector is associated with a different length of the initial interval. The possible different initial interval lengths considered are set through the 1-d grid long. Thus, $\bm{\alpha}_{t}\left[i\right]$ is the $\tilde{\alpha}$ to be used when the length of the initial interval of the observation at time $t$ is long$\left[i\right]$ (or long$\left[i\right]$ is the closest of all those considered in long). The update of $\bm{\alpha}_{t}$ is done in the same way as in ACI. However, as the conformal correction is being done as a function of interval length, only the positions associated with that interval length (and close to it) are updated. To do this, the weight vector $\bm{w}_{t}$ is constructed through a Gaussian kernel, so a new parameter $\sigma$ related to the amplitude of the kernel effect is introduced. The difference between the ACI and WACI methods throughout iterations is shown in Figure 3.
@@
The upper graph in Figure 3 shows the first iteration of both methods. Here the methods coincide, as both start from the target $\alpha$. In the next iteration (second graph from above) the methods vary, although very slightly: the ACI method uses a slightly different $\alpha$ but for all possible interval lengths, while WACI only makes this modification for initial interval lengths close to those of the previous observation. The ACI method always shows a horizontal line, while WACI shows variations, to the point of using very different alphas at “close” lengths. The third and forth graph from the top correspond to subsequent iterations.
In the case of the algorithm presented in (4), an exponential decay as a function of distance has been chosen. Of course, different weighting schemes can be considered. A scheme of fixed weights as a function of the position in the vector could also be considered, for example, one where the weights of each interval follow a geometric progression.
That is, let $\text{{long}}=\left(\text{{long}}_{\text{min}},\,\text{{long}}_{\text{min}}+h,\,\text{{long}}_{\text{min}}+2h,\,\dots,\,\text{{long}}_{\text{max}}\right)\in\mathbb{R}^{n+1}$ [^3] Here, we are considering $n+1$ points in the grid as we are interested in $n$ intervals so each position of $\bm{\alpha}_{t}$ is associated with each interval in contrast to the previous case where $\bm{\alpha}_{t}$ is associated with each extreme of the interval., then
$\bm{w}_{t}\left[j\right]=\lambda^{|i-j|},\quad|C_{\alpha,t}|\in\text{{int}}[i]\equiv\left[\text{{long}}\left[i\right],\text{{long}}\left[i+1\right]\right),$ (5)
where $i=i_{t}$ is the index of the interval length for the sample $t$.
[…]
Figure 4: (Top) Comparing the different weights schemes presented for the WACI algorithm. The exponential decay weight is shown before scaling. (Bottom) The behaviour of the two schemes can be very similar in practice.
Figure 4 shows the difference between the two proposed weighting schemes. Despite their differences, by selecting the parameters $\sigma$ and $\lambda$ in a certain way, the behaviour of both can be very similar.
Indeed, if the weighting scheme (5) is considered, asymptotic conditional coverage can be proved with respect to each of the intervals considered in the grid long.
@@
Let’s assume there exists $\delta\in\mathbb{N}$ such that $\bm{\alpha}_{t}\left[i\right]\in\left[-\delta,1+\delta\right]\;$ for all $i=1,\ldots,n$ and $t\in\mathbb{N}$. Let $i\in\left\{1,...,n\right\}$ such that there is an infinite number of $t\in\text{{int}}\left[i\right]$. If $T\longrightarrow\infty$ and the weighting scheme of (5) is considered, then
$\mathbb{P}\left(y_{T+1}\in\hat{C}^{c}_{\alpha^{*},T+1}\,|\,|C_{\alpha,T+1}|\in\text{{int}}\left[i\right]\right)=1-\alpha^{*},$
where $\alpha^{*}$ is the objective miscoverage rate and $|C_{\alpha,T+1}|$ is the length of the first interval produced at time step $T+1$.
@@
that can be decomposed based on the weight updated carried out during each iteration as
$\bm{\alpha}_{T+1}\left[i\right]-\bm{\alpha}_{1}\left[i\right]=\sum_{j=1}^{n}\sum_{t\in\text{{int}}\left[j\right]}\gamma\lambda^{|i-j|}\left(\alpha^{*}-\text{err}_{t}\right)$
Denoting by $b_{k}=\frac{\bm{\alpha}_{T+1}\left[k\right]-\bm{\alpha}_{1}\left[k\right]}{\gamma}$ and $c_{k}=\sum_{t\in\text{{int}}\left[k\right]}\left(\alpha^{*}-\text{err}_{t}\right)$ for $k=1,\dots,n$;
@@
$\underbrace{\begin{pmatrix}b_{1}\\ b_{2}\\ \vdots\\ b_{n}\\ \end{pmatrix}}_{\bm{b}}=\underbrace{\begin{pmatrix}1&\lambda^{-1}&\lambda^{-2}&\dots&\lambda^{-(n-1)}\\ \lambda^{-1}&1&\lambda^{-1}&\dots&\lambda^{-(n-2)}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ \lambda^{-(n-1)}&\lambda^{-(n-2)}&\lambda^{-(n-3)}&\dots&1\\ \end{pmatrix}}_{L}\underbrace{\begin{pmatrix}c_{1}\\ c_{2}\\ \vdots\\ c_{n}\\ \end{pmatrix}}_{\bm{c}}$
The matrix $L$ is a Toeplitz matrix equivalent to the correlation matrix of a Markov-1 signal. As discussed in Britanak et al. (2007), the inverse of $L$ exists (and it is known) and, therefore,
$\bm{c}=L^{-1}\bm{b}.$
Let $i\in\left\{1,\ldots,n\right\}$ such that $\left\{t\in\text{{int}}[i]\cap\mathbb{N}\right\}$ has an infinite number of elements and let $T_{i}=\#\left\{t\in\text{{int}}[i]:t=1,\ldots,T\right\}$. Then, as $L^{-1}$ and $\bm{b}$ are bounded, we have
$\lim_{T\rightarrow\infty}\frac{1}{T_{i}}\lVert\bm{c}\rVert_{2}=\lim_{T\rightarrow\infty}\frac{1}{T_{i}}\lVert L^{-1}\bm{b}\rVert_{2}=0.$
@@
$\lim_{T\rightarrow\infty}\frac{1}{T_{i}}\bm{c}=\bm{0}\Longrightarrow\lim_{T\rightarrow\infty}\frac{c_{i}}{T_{i}}=\lim_{T\rightarrow\infty}\frac{1}{T_{i}}\sum_{t\in\text{{int}}\left[i\right]}\text{err}_{t}-\alpha^{*}=0,$
@@
$\mathbb{P}\left(y_{T+1}\in\hat{C}^{c}_{\alpha^{*},T+1}\,|\,|C_{\alpha,T+1}|\in\text{{int}}\left[i\right]\right)\underset{T\longrightarrow\infty}{=}1-\alpha^{*}$
@@
The only assumption made to obtain the result is that the value of $\bm{\alpha}$ is bounded for every position. Although this is not formally proven, it seems a reasonable feature of the algorithm. Considering that every time the value of a certain position of $\bm{\alpha}$ is greater than 1 or less than 0, if the $\alpha$ of that position is used in a certain iteration, it is forced to reduce or increase that value respectively, which controls the explosion of that value. The only possible way that there are no limits on the value of a certain $\bm{\alpha}$ position is that once the value 1 (from above) or 0 (from below) is exceeded, that position is not worked on again too many times when compared to others and that distant positions continuously increase or decrease that value at a faster rate. Such behaviour is irrational, as one expects to go through the different $\bm{\alpha}$ positions over the iterations more or less uniformly and altering the values both up and down.
4 Real data examples

### v2

@@ 3.2 Width-Adaptive Conformal Inference
The second property we aim to achieve is maintaining the same level of confidence in the coverage level regardless of the prediction’s difficulty. In other words, the coverage should remain independent of the complexity of the situation, as is the case for the true quantiles we seek to estimate (Feldman et al., 2021). When properly estimated, the interval length serves as an indicator of prediction difficulty, reflecting both aleatoric and epistemic uncertainty. In our case, this uncertainty is captured through a quantile regression process, such as HQR. Let us denote this interval at time $T+1$ by $\widehat{C}_{\alpha,T+1}$. To ensure the desired property, we build a second interval, $\widehat{C}^{c}_{\alpha,T+1}$, by modifying $\widehat{C}_{\alpha,T+1}$. This adjusted interval is designed to satisfy property (2). From now on, the first initial interval $\widehat{C}_{\alpha,T+1}$ is called the unconformalized interval and the second one $\widehat{C}^{c}_{\alpha,T+1}$ the conformalized interval.
As we are working with time series, we will modify the ACI method to apply a different $\alpha$ as a function of time, like the original method, and also as a function of the length of the interval. Given a range of possible interval lengths from the unconformalized interval, the objective is to partition this space into smaller sub-intervals, with each sub-interval receiving a distinct correction. In other words, there is a different correction depending on the unconformalized interval length. This approach enables the differentiation of varying levels of uncertainty in the data. For instance, in Electricity Price Forecasting, two distinct regimes often emerge: one in which prices are high and the share of renewables in the energy mix is relatively small, and another in which prices are low yet uncertainty grows due to a higher proportion of renewables. The quantile regression model generating the unconformalized interval may exhibit different behaviours in these two states or may not differentiate between them when it should. In any case, it is appropriate to apply different corrections for each state. This methodology allows for the unified treatment of these states within a single framework, while also being suitable for time series data.
@@
Let $\mathcal{S}$ be the conformity scores (Section 2.3.1). Given a step $\delta\in\mathbb{R}^{+}$, the 1-d grid $\bm{L}$ is defined as $\bm{L}=\left(L_{\text{min}},L_{\text{min}}+\delta,L_{\text{min}}+2\delta,\dots,L_{\text{max}}\right)$ whose elements belong in $\mathbb{R}$. Let $\alpha^{*}$ be the objective miscoverage rate. Let’s denote the element in position $i$ of a vector $\bm{v}$ by $\bm{v}\left[i\right]$, the $p$ power of $\bm{v}$ as the $p$ power of each one of the elements of $\bm{v}$ and the absolute value of $\bm{v}$ as the absolute values of each one of the elements of $\bm{v}$. The application of WACI (Width-Adaptive Conformal Inference) over the CQR procedure looks as follows. Let $\bm{\alpha}_{1}=\left(\alpha^{*},\alpha^{*},\dots,\alpha^{*}\right)$ with the same dimension as $\bm{L},\text{err}_{1}=0,\gamma,\sigma>0.$
@@
$\begin{cases}\bm{\alpha}_{t+1}&=\bm{\alpha}_{t}+\gamma\bm{w}_{t}(\alpha^{*}-\text{err}_{t})\\ \text{err}_{t}&=\begin{cases}1&\text{if }y_{t}\not\in\widehat{C}_{\alpha^{*},t}^{c}\\ 0&\text{otherwise}\end{cases}\\ \text{{dist}}_{t}&=|\bm{L}-|\widehat{C}_{\alpha,t}||\\ \bm{w}_{t}&=\dfrac{\text{exp}\left(\frac{-\text{{dist}}_{t}^{2}}{2\sigma^{2}}\right)}{\max\left\{\text{exp}\left(\frac{-\text{{dist}}_{t}^{2}}{2\sigma^{2}}\right)\right\}}\\ \text{idx}_{t+1}&=\operatorname*{\arg\!\min}_{i}\left\{\bm{L}\left[i\right]-|\widehat{C}_{\alpha,t+1}|\right\}\\ \tilde{\alpha}_{t+1}&=\bm{\alpha}_{t+1}\left[\text{idx}_{t+1}\right]\\ \widehat{C}_{\alpha^{*},t+1}^{c}&=\left[\widehat{l}_{\alpha^{*}_{t}}^{c},\widehat{u}_{\alpha^{*}_{t}}^{c}\right]=\left[\widehat{l}_{\alpha^{*},t+1}-Q_{1-\tilde{\alpha}_{t+1}}(\mathcal{S}),\widehat{u}_{\alpha^{*},t+1}+Q_{1-\tilde{\alpha}_{t+1}}(\mathcal{S})\right]\\ \end{cases}$ (7)
@@
The first difference that can be seen with the ACI method is that in this case there is not a single $\alpha_{t}$ in each iteration, but a vector $\bm{\alpha}_{t}$. This is done in order to be able to differentiate the real scalar $\tilde{\alpha}_{t}$ that will actually be used in that iteration, which will depend on the length of the unconformalized interval. That is, each element of the vector is associated with a different length of the initial interval. The possible unconformalized interval lengths considered are set through the 1-d grid $\bm{L}$. Thus, $\bm{\alpha}_{t}\left[i\right]$ is the $\tilde{\alpha}$ to be used when the length of the initial interval of the observation at time $t$ is $\bm{L}\left[i\right]$ (or $\bm{L}\left[i\right]$ is the closest of all those considered in $\bm{L}$). The update of $\bm{\alpha}_{t}$ is done in the same way as in ACI. However, as the conformal correction is being done as a function of interval length, only the positions associated with that interval length (and close to it) are updated. To do this, the weight vector $\bm{w}_{t}$ is constructed through a Gaussian kernel, so a new parameter $\sigma$ related to the amplitude of the kernel effect is introduced. The difference between the ACI and WACI methods throughout iterations is shown in Figure 4.
The upper left graph in Figure 4 illustrates the first iteration of both methods, where they coincide as both start with the target $\alpha$. In the next iteration (upper right graph), the methods begin to diverge, although very slightly. The ACI method adjusts $\alpha$ for all possible interval lengths, whereas the WACI method only modifies $\alpha$ for unconformalized interval lengths close to those of the previous observation. The ACI method correction will always remain constant, displaying a horizontal line since it does not differentiate between interval lengths. On the contrary, WACI exhibits variations, using significantly different alphas at “close” interval lengths. The lower graphs correspond to subsequent iterations. For example, in the bottom right graph, for the next iteration, if the unconformalized interval length is around 30 or 60, the correction applied is actually bigger (in the sense that $\tilde{\alpha}>\alpha^{*}$) compared to the standard correction that CQR would use. However, if the unconformalized interval length is around 40, the correction is smaller. Such distinctions cannot be made by ACI.
3.2.1 Weighting Scheme Considerations
In the Algorithm (7), an exponential decay function of distance has been chosen for the weighting scheme. However, alternative weighting schemes could be considered. For example, a scheme with fixed weights based on the position in the vector could also be used, such as one where the weights of each interval follow a geometric progression.
@@ 3.2.1 Weighting Scheme Considerations
Figure 5: (a) Comparing the different weight schemes for the WACI algorithm. The exponential decay weight is shown before scaling. (b) The behaviour of the two schemes can be very similar in practice.
That is, let $\bm{L}=\left(L_{\text{min}},\,L_{\text{min}}+\delta,\,L_{\text{min}}+2\delta,\,\dots,\,L_{\text{max}}\right)\in\mathbb{R}^{*}$ and $\mathcal{L}(i_{t})=\left[\bm{L}\left[i_{t}\right],\bm{L}\left[i_{t}+1\right]\right)$ with $i_{t}$ the index of the interval length for the sample of instant $t$. Then,
$\bm{w}_{t}\left[j\right]=\lambda^{|i_{t}-j|},\quad|C_{\alpha,t}|\in\mathcal{L}(i_{t})$ (8)
Figure 5 shows the difference between the two proposed weighting schemes. Despite their differences, by selecting the parameters $\sigma$ and $\lambda$ in a certain way, the behaviour of both is very similar.
If the weighting scheme (8) is considered, asymptotic conditional coverage can be proved with respect to each of the intervals considered in the grid $\bf{L}$.
@@
Let’s assume there exists $\nu\in\mathbb{N}$ such that $\bm{\alpha}_{t}\left[i\right]\in\left[-\nu,1+\nu\right]\;$ for all $i=1,\ldots,n$ and $t\in\mathbb{N}$. Let $i\in\left\{1,...,n\right\}$ such that there is an infinite number of $t\in\mathcal{L}\left(i\right)$. If $T\longrightarrow\infty$ and the weighting scheme of (8) is considered, then
$\mathbb{P}\left(y_{T+1}\in\widehat{C}^{c}_{\alpha^{*},T+1}\,\left|\;\left|\widehat{C}_{\alpha^{*},T+1}\right|\right.\in\mathcal{L}\left(i\right)\right)\underset{T\longrightarrow\infty}{\longrightarrow}1-\alpha^{*},$
where $\alpha^{*}$ is the objective miscoverage rate and $|\widehat{C}_{\alpha^{*},T+1}|$ is the length of the first interval produced at time step $T+1$.
@@
Consider the set of indices of instants in whose iteration the length of the unconformalized interval belonged to the interval grid $j$. That is,
$\mathcal{I}_{j}=\left\{t\,:\,\left|\widehat{C}_{\alpha^{*},t}\right|\in\mathcal{L}\left(j\right),\,t=1,\dots,T\right\}.$
Then, the previous expression can be decomposed based on the weight updated carried out during each iteration as
$\bm{\alpha}_{T+1}\left[i\right]-\bm{\alpha}_{1}\left[i\right]=\sum_{j=1}^{n}\sum_{t\in\mathcal{I}_{j}}\gamma\lambda^{|i-j|}\left(\alpha^{*}-\text{err}_{t}\right)$
Denoting by $b_{k}=\frac{\bm{\alpha}_{T+1}\left[k\right]-\bm{\alpha}_{1}\left[k\right]}{\gamma}$ and $c_{k}=\sum_{t\in\mathcal{I}_{k}}\left(\alpha^{*}-\text{err}_{t}\right)$ for $k=1,\dots,n$;
@@
$\underbrace{\begin{pmatrix}b_{1}\\ b_{2}\\ \vdots\\ b_{n}\\ \end{pmatrix}}_{\bm{b}}=\underbrace{\begin{pmatrix}1&\lambda^{-1}&\lambda^{-2}&\dots&\lambda^{-(n-1)}\\ \lambda^{-1}&1&\lambda^{-1}&\dots&\lambda^{-(n-2)}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ \lambda^{-(n-1)}&\lambda^{-(n-2)}&\lambda^{-(n-3)}&\dots&1\\ \end{pmatrix}}_{\Lambda}\underbrace{\begin{pmatrix}c_{1}\\ c_{2}\\ \vdots\\ c_{n}\\ \end{pmatrix}}_{\bm{c}}$
The matrix $\Lambda$ is a Toeplitz matrix equivalent to the correlation matrix of a Markov-1 signal. As discussed in Britanak et al. (2007), the inverse of $\Lambda$ exists (and it is known) and, therefore,
$\bm{c}=\Lambda^{-1}\bm{b}.$
Let $i\in\left\{1,\ldots,n\right\}$ such that $\left\{t\in\mathcal{I}_{i}\cap\mathbb{N}\right\}$ has an infinite number of elements and let $T_{i}=\#\left\{t\in\mathcal{I}_{i}:t=1,\ldots,T\right\}$. Then, as $\Lambda^{-1}$ and $\bm{b}$ are bounded, we have
$\lim_{T\rightarrow\infty}\frac{1}{T_{i}}\lVert\bm{c}\rVert_{2}=\lim_{T\rightarrow\infty}\frac{1}{T_{i}}\lVert\Lambda^{-1}\bm{b}\rVert_{2}=0.$
@@
$\lim_{T\rightarrow\infty}\frac{1}{T_{i}}\bm{c}=\bm{0}\Longrightarrow\lim_{T\rightarrow\infty}\frac{c_{i}}{T_{i}}=\lim_{T\rightarrow\infty}\frac{1}{T_{i}}\sum_{t\in\mathcal{I}_{i}}\text{err}_{t}-\alpha^{*}=0,$
@@
$\mathbb{P}\left(y_{T+1}\in\widehat{C}^{c}_{\alpha^{*},T+1}\,\left|\;\left|\widehat{C}_{\alpha^{*},T+1}\right|\right.\in\mathcal{L}(i)\right)\underset{T\longrightarrow\infty}{\longrightarrow}1-\alpha^{*}$
@@
Although this is not exactly the condition represented by equation (2), coverage is still achieved depending on the complexity of the forecast. Therefore, it is important that the unconformalized interval shows the desired relationship between interval length and prediction difficulty (which it is achieved by applying HQR in the first step).
@@
The only assumption made to obtain the result is that the value of $\bm{\alpha}$ is bounded for every position. Although this is not formally proven, it seems a reasonable feature of the algorithm. If the value of a certain position of $\bm{\alpha}$ exceeds 1 or falls below 0 during any iterations, it is forced to decrease or increase accordingly, thereby controlling the explosion of that value. The only scenario where no limits exists on a certain position of $\bm{\alpha}$ is when, after surpassing 1 (from above) or 0 (from below), that position is no longer frequently updated compared to others. In such cases, distant positions might continuously increase or decrease at a faster rate. This behaviour is irrational, as one would expect the algorithm to update different $\bm{\alpha}$ positions uniformly over iterations, adjusting the values both upward and downward.
Another implicit assumption is that $\delta$ is sufficiently small, meaning the distance between grid separator points is not excessively large. This is important because, if the grid intervals are too wide, a scenario may arise where all initial intervals fall within the same grid interval. In such cases, the procedure essentially reduces to applying ACI. When using a distance-based weighting scheme (as in equation (7)), choosing a very fine grid generally does not present a problem, aside from potentially increasing the algorithm’s computational runtime. However, for position-based weighting schemes (as in equation (8)), the decay of the weights along positions must reasonably reflect the distances between observations. Specifically, the weight decay should adjust based on the choice of $\delta$: if $\delta$ is small relative to the scale of the problem, the weight decay should be smoother, whereas a larger $\delta$ requires a steeper weight decay. In any case, it is recommended using as fine a grid as possible to ensure better granularity and accuracy in the results.
We emphasize that while HQR and WACI can be applied independently, combining them is essential to ensure that the desired properties are achieved.
4 Computational experiments